If f(1) = -3 and f(x+2) = 3*f(x) - 3, find f(9).
Answer by jim_thompson5910(35256)

f(1) = -3

f(x+2) = 3*f(x) - 3

f(1+2) = 3*f(1) - 3

f(1+2) = 3*(-3) - 3

f(3) = -9 - 3

f(3) = -12

-------------------------------------------------------

f(x+2) = 3*f(x) - 3

f(3+2) = 3*f(3) - 3

f(5) = 3*(-12) - 3

f(5) = -36 - 3

f(5) = -39

-------------------------------------------------------

f(x+2) = 3*f(x) - 3

f(5+2) = 3*f(5) - 3

f(7) = 3*(-39) - 3

f(7) = -117 - 3

f(7) = -120

-------------------------------------------------------

f(x+2) = 3*f(x) - 3

f(7+2) = 3*f(7) - 3

f(9) = 3*(-120) - 3

f(9) = -360 - 3

f(9) = -363

So the answer is f(9) = -363